National Corporation needs to set a target price for its newly designed product M14–M16. The following data relate to this new product.

Per Unit Total
Direct materials $30
Direct labor $44
Variable manufacturing overhead $11
Fixed manufacturing overhead $1,264,000
Variable selling and administrative expenses $ 5
Fixed selling and administrative expenses $ 1,106,000


These costs are based on a budgeted volume of 79,000 units produced and sold each year. National uses cost-plus pricing methods to set its target selling price. The markup percentage on total unit cost is 40%.

Compute the total variable cost per unit, total fixed cost per unit, and total cost per unit for M14–M16.

Variable cost per unit $enter a dollar amount
Fixed cost per unit enter a dollar amount
Total cost per unit

$enter a total of the two previous amounts

Compute the desired ROI per unit for M14–M16.

Desired ROI $enter the desired ROI per unit in dollars per unit

Compute the target selling price for M14–M16.

Target selling price per unit $enter the target selling price per unit in dollars

Compute variable cost per unit, fixed cost per unit, and total cost per unit assuming that 59,250 M14–M16s are produced and sold during the year.

Variable cost per unit $enter a dollar amount
Fixed cost per unit enter a dollar amount
Total cost per unit $enter a total of the two previous amounts
Answer #1

Variable cost per unit = Direct material + Direct labor + Variable manufacturing overhead + Variable selling and admin expense

= 30+44+11+5

= $90 per unit

Fixed cost per unit = Total cost/Number of units
= (1,264,000+1,106,000)/79000

= $30 per unit

Total cost per unit = 90+30 = $120 per unit

Desired ROI = 120*40% = $48 per unit

Target Selling price = 120+48 = $168 per unit

When 59,250 units are sold:

Variable cost per unit = $90

Fixed cost per unit = (1,264,000+1,106,000)/59,250 = $40 per unit

Total cost per unit = $130 per unit